(fog)(x)(gof), find f(x)=(3)/(1-8x), g(x)= (1)/(x)
Answer by stanbon(75887)

(fog)(x)(gof), find f(x)=(3)/(1-8x), g(x)= (1)/(x)
---------------
fog(x) = f[1/x] = 3/(1-8/x) = 3x(x-8)
-----------------
gof(x) = g[3/(1-8x)] = (1-8x)/3
-------
Ans:: (3x)(x-8)(1-8x)/3 = x(x-8)(1-8x) = (x^2-8x)(1-8x) = x^2+64x^2-8x^3-8x
-----
= -8x^3+65x^2-8x
-----------------
Cheers,
Stan H.
----------------------------